# Fast Feature Matching of UAV Images via Matrix Band Reduction-based GPU Data Schedule

### v1

@@ 2.2.3 Cascade hashing matching enhanced by geometric constraints
For each block row $\{{block}_{i1},{block}_{i2},\ldots,{block}_{il}\}$, the above pipeline is executed to obtain initial candidate matches for all related match pairs. Noticeably, in step (1), redundant data loading operations can be avoided. As illustrated in Figure 3(b) for processing the second block row, only descriptors in the last block ${block}_{il}$ rendered in orange color should be loaded into GPU since all the other required data has been loaded when processing the first block row. Besides, after processing each block row, data in the first block ${block}_{i1}$ rendered in green color is free as in step (3). The interleave free-load operation goes through the whole pipeline.
False matches inevitably existed in initial matches due to limited discriminability of local feature descriptors and large perspective deformations. In this study, outlier removal is finally conducted by combining the spatial angular order (SAO) based local geometric constraint and RANSAC-based global geometric constraint (Jiang et al., 2020b). The local geometric constraint is designed by using the k-nearest neighbors deduced from the Delaunay triangulation and its corresponding graph. To achieve further efficiency improvement, outlier removal is simultaneously executed in CPU after obtaining initial matches in step (2). Thus, the results retained in outlier removal are the final matches.
@@ 2.3 Algorithm implementation
According to the workflow shown in Figure 1, this study implements the proposed cascade hashing matching algorithm by using the C++ programming language. For match pair retrieval, SIFTGPU (Wu, 2011) with default parameters is utilized to extract SFIT features, and Lloyd’s K-means cluster algorithm (Lloyd, 1982) is adopted to cluster selected training descriptors for the codebook construction. Besides, with the aid of VLFeat (Vedaldi and Fulkerson, 2010) and FAISS packages (Johnson et al., 2019), we have implemented the VLAD-HNSW based image retrieval solution. To facilitate performance evaluation in SfM reconstruction, the cascade hashing feature matching solution has been integrated into the SfM engine in Jiang et al. (2022b). The pseudo-code is presented in Algorithm 1.
@@
3:     Select a subset of $p$ percent of images from $I$ as training images $I_{p}$
@@
10:              Load descriptors in current block ${block}_{ij}$
11:              Execute cascade hashing matching for ${block}_{ij}$
12:              Remove outliers and store results $match\leftarrow\left\{m_{ij}\right\}$
13:              Free unnecessary data in block ${block}_{i1}$
14:              Set matrix $M_{sym}$ cells as zero for processed pairs
@@ 3.1 Datasets and software packages
Three UAV datasets are collected in study for performance evaluation. The details of data acquisitions and collected images are presented in Table 1 and described as follows.
@@
    The second dataset encompasses 4,030 images from a complex university building as shown in Figure 4(b). These images are captured by a DJI M300 RTK UAV, equipped with a DJI Zenmuse P1 camera that produces images with dimensions of 8,192 by 5,460 pixels. The dataset has been collected by using optimized views photogrammetry (Li et al., 2023), which adjusts camera orientation according to geometric attributes of terrestrial objects. The GSD is about 1.2 cm. For absolute orientation, 26 Ground Control Points (GCPs) have been gathered using a total station, which offers a nominal precision of 0.8 cm and 1.5 cm in the horizontal and vertical direction, respectively.
@@
    The third dataset is captured from a test site that is predominantly characterized by low-rise structures and an abundance of dense vegetation, with a river flowing through the area as shown in Figure 4(c). For data acquisition, a typical penta-view oblique photogrammetry with five SONY ILCE 7R cameras has been adopted, which record images with dimensions of 6,000 by 4,000 pixels. At an altitude of 87.1 m, a collection of 21,654 images is gathered, achieving a GSD of 1.2 cm.
Table 1: Detailed information about the three UAV datasets.
@@
Method Match pair selection Feature matching algorithm Data schedule Hardware acceleration Software version ColMap-CPU Input pairs KD-Tree No Multi-core CPU 3.10 ColMap-GPU Input pairs NNS Yes GPU & CPU 3.10 AliceVision Input pairs Cascade hashing No Multi-core CPU v2023.3.0 TLBDS Input pairs Cascade hashing Yes GPU & CPU / Metashape POS / Yes GPU & CPU 1.8.21 Pix4Dmapper VOC-Tree / Yes GPU & CPU 4.4.12 Ours Input pairs Cascade hashing Yes GPU & CPU /
For performance evaluation, this study compares the proposed algorithm with five well-known software packages, as listed in Table 2, including two open-source software packages ColMap (Schonberger and Frahm, 2016) and AliceVision (Griwodz et al., 2021), two commercial software packages Metashape (Agisoft, 2024) and Pix4Dmapper (Pix4Dmapper, 2024), and one recently cascade hashing matching algorithm, termed TLBDS (Zhang et al., 2023). In the experiments, all tests are conducted on a Windows platform equipped with 64 GB memory, a 3.0 GHz Intel Core i9-13900K CPU and a 24 GB NVIDIA GeForce RTX 4090 graphic card.
3.2 Analysis of schedule block size and ratio test threshold on feature matching
In the proposed cascade hashing based feature matching algorithm, the schedule block size ${Size}_{blk}$ and ratio test threshold $t_{r}$ are two critical parameters. The schedule block size determines the quantity of data that is loaded and free while processing each block row and the count of match pairs that are fed to GPU to perform feature matching. For a small block size, the load and free of data are consistently less demanding, which in turn results in reduced GPU usage. Conversely, a large block size can process more matching pairs simultaneously, thereby increasing the GPU usage and data scheduling demands. Meanwhile, the ratio test threshold determines the number of initial matches that are retained in cascade hashing feature matching. In contrast to global NNS, candidate matches have been filtered in the coarse matching step, which leads to the practical threshold used in the other algorithm is not proper. Thus, we will analysis their influence on feature matching and select their optimal values.
For the analysis of schedule block size ${Size}_{blk}$, datasets 1 and 2 are selected as they are captured by using two different photogrammetric configurations, which results in the collected images with varying viewing directions and image overlap degrees. In this test, the schedule block size is sampled with the values of 50, 100, 150, 200, 400, 600, and 800 under the limitation of the maximum GPU memory, and time cost is used as the metric for performance evaluation. The results are shown in Figure 5. It is clearly shown that the time cost decreases consistently with the increase of schedule block size ranging from 50 to 400. The main reason is that for the large block size, more match pairs are fed to GPU at one time, which would increase the GPU usage and in turn increase matching efficiency. Noticeably, higher efficiency improvement is observed from dataset 2 when compared with that in dataset 1. It can be explained by the larger image overlap degree in dataset 2. When the schedule block size further increases from 400 to 800, it is shown that the time cost almost keeps constant for these two datasets because of the full usage of the maximum GPU computing power. Therefore, ${Size}_{blk}$ is configured as 400 in the following tests.
@@ 3.2 Analysis of schedule block size and ratio test threshold on feature matching
3.3 Analysis of the key steps of the proposed feature matching workflow
@@
To verify the effectiveness of the proposed data schedule strategy, this study compares feature matching efficiency without and with the data schedule. The only difference between these two tests is the usage of the data schedule strategy. The performance is evaluated by the metrics time and speed, which are quantified respectively by time costs of feature matching and the number of pairs processed per second. Table 4 presents the results. It is shown that when not using the data schedule strategy, the matching speed is merely 18.94 pairs/s, 23.01 pairs/s, and 23.26 pairs/s, respectively; on the contrary, by using the data schedule strategy, the matching speed improves obviously, which reaches to 340.85 pairs/s, 362.11 pairs/s, and 228.95 pairs/s for each dataset. In other words, the speedup ratio is 17.99, 15.74, and 9.84 for the three datasets, respectively. It can be explained by frequent IO burdens in data load and the low GPU usage for sequential feature matching. Thus, the proposed data schedule strategy improves the efficiency of cascade hashing feature matching.
@@
Dataset No schedule Schedule Speedup Time (s) Speed Time (s) Speed 1 3260.22 18.94 181.20 340.85 17.99 2 2946.06 23.01 187.20 362.11 15.74 3 15766.86 23.26 1602.06 228.95 9.84
3.4 Comparison with state-of-the-art methods
In this section, the performance of the proposed algorithm is compared with state-of-the-art methods and evaluated in terms of feature matching and SfM-based 3D reconstruction. The former is conducted to quantify the performance of time cost, inlier number, and inlier ratio in feature matching; the latter is utilized to verify the effectiveness in 3D reconstruction.
3.4.1 Feature matching
Feature matching is first evaluated by using four metrics, i.e., time cost, matching speed, inlier number, and inlier ratio. The metric time cost includes the time consumption in feature matching and outlier removal, and the metric matching speed is calculated as the ratio between match pairs and time costs; the metrics inlier number and inlier ratio represent the average true matches and the average number ratio between inliers and initial matches over all match pairs, respectively. In this test, six algorithms are used for performance evaluation, i.e., ColMap-GPU, ColMap-CPU, AliceVision, TLBDS, Metashape, and Pix4Dmapper. For unbiased comparison, the maximum number of matches is configured as 8,192 in Metashape and Pix4Dmapper as that used in the other algorithms.
@@ 3.4.1 Feature matching
Method Time cost (min) Matching speed (pairs/s) Dataset 1 Dataset 2 Dataset 3 Dataset 1 Dataset 2 Dataset 3 ColMap-CPU 258.4 309.8 2064.9 4.0 3.6 3.0 ColMap-GPU 3.4 3.6 71.7 302.4 315.9 85.2 AliceVision 22.7 24.6 155.3 61.3 58.5 55.8 TLBDS 11.3 12.1 70.8 91.1 93.7 86.4 Metashape 8.3 9.2 57.0 33.0 38.9 20.3 Pix4Dmapper 21.0 10.9 136.7 41.0 54.6 30.2 Ours 3.0 3.1 26.7 340.8 362.1 229.0
Figure 10: The feature matching speed of all evaluated methods.
For the further analysis, we can see that: (1) the time cost of ColMap-GPU is approximately one third of that in TLBDS for datasets 1 and 2, while they have almost the same time cost in dataset 3. The main reason is due to their different data schedule strategies. In TLBDS, images are ordered according to their connections and are sequentially fed into GPU during feature matching. It causes a lower usage ratio of GPU. On the contrary, ColMap-GPU divides match pairs into blocks that are fed into GPU at one time, which increases overall usage ratio of GPU. However, as the data schedule strategy in ColMap-GPU does not consider image connections, redundant data IO becomes serious for the large-scale dataset 3; (2) compared with TLBDS and ColMap-GPU, the proposed algorithm achieves the highest efficiency for the three datasets, whose time cost is 3.0 mins, 3.1 mins, and 26.7 mins, respectively. The main reason can be explained by the MBR-based data schedule strategy, which considers the usage ratio of GPU to increase efficiency and image connections to decrease IO burden.
@@
Figure 11 presents the comparison in terms of inlier number and inlier ratio. Noticeably, these two metrics are not calculated for AliceVision, Metashape, and Pix4Dmapper since only the executable application is utilized in this test. It is shown that for the metric inlier number, ColMap-GPU, TLBDS, and the proposed algorithm achieve comparable performance, which is relatively lower than that of ColMap-CPU. For the metric inlier ratio, ColMap-GPU achieves the highest performance, which is 0.43, 0.59, and 0.60 for the three datasets. In addition, Figure 12 presents the individual plot of the inlier ratio for the three datasets. Since TLBDS uses the same cascade hashing for feature matching, only the result of the proposed algorithm is analyzed. We can see that the proposed algorithm achieves better performance compared with ColMap-CPU, and evaluated methods have almost consistent change in inlier ratios for all match pairs. In conclusion, the proposed algorithm achieves speedup ratios ranging from 77.0 to 100.0 for feature matching compared with traditional KD-Tree based matching methods and provides comparable matching results for subsequent SfM reconstruction.
3.4.2 SfM-based 3D reconstruction
SfM-based 3D reconstruction is then executed by using the obtained feature matches. In general, the performance of SfM-based 3D reconstruction is affected by the inlier number and correspondence distribution over image planes, which can be utilized to evaluate the quality of feature matching. SfM-based 3D reconstruction can be executed without GCPs for relative orientation and with GCP for absolute orientation. For relative BA without GCPs, SfM is used to compare the efficiency and completeness of reconstructed models, in which the efficiency is quantified by time costs, and the completeness is measured by the number of registered images and 3D points. In this test, the parallel SfM solution proposed in Jiang et al. (2024) is used as the engine for ColMap-GPU, TLBDS, and the proposed algorithm.

### v2

@@ 2.3.2 Schedule block guided cascade hashing feature matching
For each block row $\{{block}_{i1},{block}_{i2},\ldots,{block}_{il}\}$, the above pipeline is executed to obtain initial candidate matches for all involved match pairs. Noticeably, in step (1), redundant data loading operations can be avoided. As illustrated in Figure 5(b) for processing the second block row, only descriptors in the last block ${block}_{il}$ rendered in orange color should be loaded into GPU since all the other required data has been loaded when processing the first block row. Besides, after processing each block row, data in the first block ${block}_{i1}$ rendered in green color is free as in step (3). The interleave free-load operation goes through the whole pipeline.
2.3.3 Outlier removal combining local and global geometric constraints
False matches inevitably existed in initial matches due to limited discriminability of local feature descriptors and large perspective deformations. In this study, outlier removal is finally conducted by combining the spatial angular order (SAO) based local geometric constraint and RANSAC-based global geometric constraint (Jiang et al., 2020b). The local geometric constraint is designed by using the k-nearest neighbors deduced from the Delaunay triangulation and its corresponding graph, as shown in Figure 7. For one candidate match $p_{i},q_{i}$}, their K nearest neighbors $p_{N_{i}}$ and $q_{N_{i}}$ are first found and ordered based on their angular order with respect to the horizonal axis. For example, the K nearest neighbors of $p_{i}$ are $O_{p_{i}}={{p_{N_{5}},p_{N_{4}},p_{N_{3}},p_{N_{2}},p_{N_{1}},p_{N_{6}}}}$. Similarly, the K nearest neighbors $O_{q_{i}}$ of $q_{i}$ can also be found. The dissimilarity score $Score$ of the candidate match $p_{i},q_{i}$} is then calcualted by using the cyclic edit distance (CED) algorithm between these two sets of nearest neighbors, i.e., $Score=\frac{1}{N}d_{ced}(O_{p_{i}},O_{q_{i}})$. $N$ is the number of neighbors; $d_{ced}(.)$ calculates the CED. To achieve further efficiency improvement, outlier removal is simultaneously executed in CPU after obtaining initial matches in step (2), in which initial matches with larger dissimilarity scores are prone to false matches. Thus, the results retained in outlier removal are the final matches.
Figure 7: The principle of the SAO-based local geometric constraint.
@@ 2.4 Algorithm implementation
According to the workflow presented in Figure 2, this study implements the proposed feature matching algorithm by using the C++ programming language. For match pair retrieval, SIFTGPU (Wu, 2011) with default parameters is utilized to extract SFIT features, and Lloyd’s K-means cluster algorithm (Lloyd, 1982) is adopted to cluster selected training descriptors for the codebook construction. With the aid of VLFeat (Vedaldi and Fulkerson, 2010) and FAISS packages (Johnson et al., 2019), we have implemented the VLAD-HNSW based image retrieval solution. To facilitate performance evaluation in SfM reconstruction, the cascade hashing feature matching solution has been integrated into the SfM engine in Jiang et al. (2022b). The pseudo-code is presented in Algorithm 1. Noticeable, in the procedure CascadeHashingMatching, lines 4 to 8 for MBR-based schedule block generation are executed in CPU, and lines 10 to 14 uses both GPU and CPU for feature matching and outlier removal. These two steps operate between CPU and GPU. The symbols $\uparrow$ and $\downarrow$ indicate uploading data from host memory to GPU memory and releasing data in GPU, respectively. The symbols $\equiv$ and - indicate feature matching in GPU and outlier removal in CPU, respectively. These fives steps are executed asynchronously.
@@
3:  Select a subset of $p$ percent of images from II as training images $I_{p}$
@@
10:     $\uparrow$ Load descriptors in current block ${block}_{ij}$
11:     $\equiv$ Execute cascade hashing matching for ${block}_{ij}$
12:     - Remove outliers and store results $match\leftarrow\left\{m_{ij}\right\}$
13:     $\downarrow$ Free unnecessary data in block ${block}_{i1}$
14:     - Set matrix $M_{sym}$ cells as zero for processed pairs
@@ 3 Experimental results
Three large-scale UAV datasets are utilized for performance evaluation. First, we analyze the influence of the ratio test threshold and schedule block size on feature matching precision and efficiency. Second, the key steps of the proposed solution are evaluated using ablation studies of feaure matching. Finally, we compare the proposed algorithm with open-source and commercial software packages in terms of feature matching and SfM-based 3D reconstruction. In this study, all algorithms have been implemented by using the C++ programming language, and the tests are conducted on a Windows platform equipped with 64 GB memory, a 3.0 GHz Intel Core i9-13900K CPU and a 24 GB NVIDIA GeForce RTX 4090 graphic card.
@@ 3.1 Datasets and software packages
Three UAV datasets have been collected in study for performance evaluation. The details of data acquisitions are presented in Table 1.
Table 1: Detailed information about the three UAV datasets.
@@
    The second dataset encompasses 4,030 images from a complex university building as shown in Figure 8(b). These images are captured by a DJI M300 RTK UAV, equipped with a DJI Zenmuse P1 camera that produces images with dimensions of 8,192 by 5,460 pixels. The dataset has been collected using optimized views photogrammetry (Li et al., 2023), which adjusts camera orientation according to geometric attributes of terrestrial objects. The GSD is about 1.2 cm. For absolute orientation, 26 Ground Control Points (GCPs) have been gathered using a total station, which offers a nominal precision of 0.8 cm and 1.5 cm in the horizontal and vertical direction, respectively.
@@
    The third dataset is captured from a test site that is characterized by low-rise structures and an abundance of dense vegetation, with a river flowing through the area as shown in Figure 8(c). For data acquisition, a typical penta-view oblique photogrammetry with five SONY ILCE 7R cameras has been adopted, which record images with dimensions of 6,000 by 4,000 pixels. At an altitude of 87.1 m, a collection of 21,654 images is gathered, achieving a GSD of 1.2 cm.
@@
Method Match pair selection Feature matching algorithm Data schedule Hardware acceleration Software version ColMap-CPU Input pairs KD-Tree No Multi-core CPU 3.10 ColMap-GPU Input pairs NNS Yes GPU & CPU 3.10 AliceVision Input pairs Cascade hashing No Multi-core CPU v2023.3.0 TLBDS Input pairs Cascade hashing Yes GPU & CPU / GraphPart Input pairs Cascade hashing Yes GPU & CPU / Metashape POS / Yes GPU & CPU 1.8.21 Pix4Dmapper VOC-Tree / Yes GPU & CPU 4.4.12 Ours Input pairs Cascade hashing Yes GPU & CPU /
For performance evaluation, this study compares the proposed algorithm with six well-known solutions, as presented in Table 2, including two open-source software packages ColMap (Schonberger and Frahm, 2016) and AliceVision (Griwodz et al., 2021), two commercial software packages Metashape (Agisoft, 2024) and Pix4Dmapper (Pix4Dmapper, 2024), one classical graph partition based matching algorithm GraphPart, and one recently cascade hashing matching algorithm TLBDS (Zhang et al., 2023). GraphPart is implmented by splitting the whole view graph into sub-graphs with pre-defined vertex number.
3.2 Analysis of schedule block size and ratio test threshold
In the proposed feature matching algorithm, the schedule block size ${Size}_{blk}$ and ratio test threshold $t_{r}$ are two critical parameters. The schedule block size ${Size}_{blk}$ determines the data quantity that is loaded while processing each block row and the count of match pairs that are fed to GPU. For a small block size, the load and free of data are consistently less demanding, which in turn results in reduced GPU usage. Conversely, a large block size can process more matching pairs simultaneously, thereby increasing GPU usage and data scheduling demands. The ratio test threshold $t_{r}$ determines the number of initial matches that are retained in the cascade hashing feature matching. In contrast to global NNS, candidate matches have been filtered in the coarse matching step, which leads to the practical threshold used in the other algorithm being improper. Thus, we will analyze their influence on feature matching and select their optimal values.
@@ 3.2 Analysis of schedule block size and ratio test threshold
For the analysis of schedule block size ${Size}_{blk}$, datasets 1 and 2 are selected as they are captured by using two different photogrammetric configurations, which results in the collected images with varying viewing directions and image overlap degrees. In this test, the schedule block size is sampled with the values of 50, 100, 150, 200, 400, 600, and 800 under the limitation of the maximum GPU memory, and time cost is used as the metric for performance evaluation. The results are shown in Figure 9. It is shown that the time cost decreases consistently with the increase of schedule block size ranging from 50 to 400. The main reason is that for the large block size, more match pairs are fed to GPU at one time, which would increase the GPU usage and in turn increase matching efficiency. Noticeably, higher efficiency improvement is observed from dataset 2 when compared with that in dataset 1. It can be explained by the larger image overlap degree in dataset 2. When the schedule block size further increases from 400 to 800, it is shown that the time cost almost keeps constant for these two datasets due to the full usage of the GPU computing power. Thus, ${Size}_{blk}$ is set as 400 in following tests.
@@
3.3 Analysis of the key steps of the proposed feature matching solution
3.3.1 Image retreival-based match pair selection
@@ 3.3.2 Schedule block generation based on MBR
To verify the effectiveness of the proposed data schedule strategy, this study compares feature matching efficiency without and with the data schedule. The only difference between these two tests is the usage of the MBR-based data schedule and SAO-based outlier removal. The performance is evaluated by the metrics time and speed, which are quantified respectively by time costs of feature matching and the number of pairs processed per second. Table 4 presents the results. It is shown that when not using the data schedule strategy, the matching speed is merely 18.94 pairs/s, 23.01 pairs/s, and 23.26 pairs/s, respectively; on the contrary, by using the data schedule strategy, the matching speed improves obviously, which reaches to 340.85 pairs/s, 362.11 pairs/s, and 228.95 pairs/s for each dataset. In other words, the speedup ratio is 18.0, 15.7, and 9.8 for the three datasets, respectively. It can be explained by the low GPU usage for sequential feature matching. Thus, the proposed data schedule strategy improves the efficiency of cascade hashing feature matching.
@@
Dataset No schedule Schedule Speedup Time (s) Speed Time (s) Speed 1 3260.22 18.9 181.20 340.8 18.0 2 2946.06 23.0 187.20 362.1 15.7 3 15766.86 23.3 1602.06 229.0 9.8
3.3.3 Ablation study
The ablation study is conducted for the further evaluation of each key step. In this test, TLBDS is used as the benchmark algorithm because it has almost the same key steps as the proposed solution for feature matching, including data schedule and cascade hashing. Combined with MBR based data schedule and SAO based outlier removal, two versions, i.e., Ours-MBR and Ours-Full, are evaluated. The statistical results are listed in Table 5. It is clearly shown that by using MBR, the feature matching speed of Ours-MBR increases to 235.3 pairs/s, 213.6 pairs/s, and 70.4 pairs/s compared with TLBDS, whose speedup is 4.8, 3.2, and 1.2, respectively, for the three datasets. By using both MBR and SAO, Ours-Full can achieve further acceleration in feature matching, whose speedup is 6.9, 5.5, and 4.0, respectively. We can see that the performance of proposed algorithm can benefit obviously from the usage of these two strategies in feature matching.
Table 5: The ablation study of the proposed workflow. The term D1, D2 and D3 represents dataset 1, dataset 2 and dataset 3, respectively.
Method Key step Speed (pairs/s) Speedup MBR SAO D1 D2 D3 D1 D2 D3 TLBDS $\times$ $\times$ 49.4 66.4 56.7 - - - Ours-MBR $\times$ 235.3 213.6 70.4 4.8 3.2 1.2 Ours-Full 340.8 362.1 229.0 6.9 5.5 4.0
3.4 Feature matching comparison with state-of-the-art methods
In this section, the performance of the proposed algorithm is compared with state-of-the-art methods in feaure matching. This test provides the insight comparison of the proposed solution and their fearure matching results.
3.4.1 The evaluation of feature matching efficiency
Feature matching efficiency is first evaluated by using two metrics, i.e., time cost and matching speed. The metric time cost includes the time consumption in feature matching and outlier removal, and the metric matching speed is calculated as the ratio between match pairs and time costs. In this test, seven algorithms are compared, i.e., ColMap-CPU, ColMap-GPU, AliceVision, TLBDS, GraphPart, Metashape, and Pix4Dmapper. For unbiased comparison, the maximum number of matches is configured as 8,192 in Metashape and Pix4Dmapper as that used in the other algorithms.
Method Time cost (min) Matching speed (pairs/s) Dataset 1 Dataset 2 Dataset 3 Dataset 1 Dataset 2 Dataset 3 ColMap-CPU 258.4 309.8 2064.9 4.0 3.6 3.0 ColMap-GPU 3.4 3.6 71.7 302.4 315.9 85.2 AliceVision 22.7 24.6 155.3 61.3 58.5 55.8 TLBDS 11.3 12.1 70.8 91.1 93.7 86.4 GraphPart 8.3 10.4 54.6 124.0 108.6 112.0 Metashape 8.3 9.2 57.0 33.0 38.9 20.3 Pix4Dmapper 21.0 10.9 136.7 41.0 54.6 30.2 Ours 3.0 3.1 26.7 340.8 362.1 229.0
Figure 14: The feature matching speed of all evaluated methods.
@@ 3.4.1 The evaluation of feature matching efficiency
For the further analysis, we can see that: (1) the time cost of ColMap-GPU is approximately one third of that in TLBDS for datasets 1 and 2, while they have almost the same time cost in dataset 3. The main reason is due to their different data schedule strategies. In TLBDS, images are ordered according to their connections and are sequentially fed into GPU during feature matching. Although it has a higher usage ratio of GPU memory, the weak connection between images causes a lower usage ratio of GPU computing, as shown in Figure 15(b). On the contrary, ColMap-GPU divides match pairs into blocks that are fed into GPU at one time, which increases overall usage ratio of GPU computing. However, as the data schedule strategy in ColMap-GPU does not consider image connections, the low usage ratio of GPU memory can be observed in Figure 15(a); (2) compared with ColMap-GPU, TLBDS, and GraphPart, the proposed algorithm achieves the highest efficiency for the three datasets, whose time cost is 3.0 mins, 3.1 mins, and 26.7 mins, respectively. The main reason can be explained by the MBR-based data schedule, which considers the usage ratio of GPU computing and memory to increase feature matching efficiency, as presented in Figure 15(d).
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 15: The comparison of the GPU computing and memory usage of dataset 1: (a) ColMap-GPU; (b) TLDBS; (c) GraphPart; (d) MBR. The values in the rectangle are the statitics of mean and max, respectively, for each evaluated method.
@@
3.4.2 The evaluation of feature matching resutls
Feature matching results are then evaluated by using two metrics, i.e., inlier number and inlier ratio, which represents the average true matches and the average number ratio between inliers and initial matches over all match pairs, respectively. Figure 16 presents the comparison. Noticeably, these two metrics are not calculated for AliceVision, Metashape, and Pix4Dmapper since only the executable application is utilized in this test. It is shown that for the metric inlier number, ColMap-GPU, TLBDS, and the proposed algorithm achieve comparable performance, which is relatively lower than that of ColMap-CPU. The main reason can be the usage of GPU with the single-precision floating-point format for feature matching when compared with CPU with the double-precision floating-point format. For the metric inlier ratio, ColMap-GPU achieves the highest performance, which is 0.43, 0.59, and 0.60 for the three datasets. From the setting in Table 2, it can be explained by the used feature matching algorithm. In contrast to ColMap-CPU, TLBDS and the proposed algorithm with ANN for feature matching, ColMap-GPU uses NNS for feature matching, which has higher precision generally. In addition, Figure 17 presents the individual plot of the inlier ratio for the three datasets. Since TLBDS uses the same cascade hashing for feature matching, only the result of the proposed algorithm is analyzed. We can see that the proposed algorithm achieves better performance compared with ColMap-CPU, and evaluated methods have almost consistent change in inlier ratios for all match pairs. In conclusion, the proposed algorithm achieves speedup ratios ranging from 77.0 to 100.0 for feature matching compared with traditional KD-Tree based matching methods and provides comparable matching results for subsequent SfM reconstruction.
3.5 SfM-based 3D reconstruction
SfM-based 3D reconstruction is then executed using feature matching results. In general, the performance of SfM-based 3D reconstruction is affected by the inlier number and correspondence distribution over image planes, which can be utilized to evaluate the quality of feature matching. SfM-based 3D reconstruction can be executed without GCPs for relative orientation and with GCP for absolute orientation. The details are presented as follows.
3.5.1 Relative BA without GCPs